OpenStudy (anonymous):

rationalize the denominator, then simplify: 8/ 3sqaure root of 2

OpenStudy (anonymous):

is this \[\frac{8}{\sqrt[3]{2}}\]?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

if so, you may want to say "cube root of 2" to make it clearer.

OpenStudy (anonymous):

ohh okay, sorry i forgot that

OpenStudy (anonymous):

No that can't be right. I think it's square root. Not cubic root.

OpenStudy (anonymous):

Oh okay. I guess it is.

OpenStudy (anonymous):

idk?

OpenStudy (anonymous):

ok then you want the denominator to be a perfect cube, so multiply by \[\frac{\sqrt[3]{4}}{\sqrt[3]{4}}\] because then you will get the cube root of 8 in the denominator, and that is 2

OpenStudy (anonymous):

Is it \[3 * \sqrt{2}\] or is it \[\sqrt[3]{2}\] The first means "3 times the number which, when multiplied by itself, gives 2" The second means "the number that, when cubed, gives 2."

OpenStudy (anonymous):

so your answer is \[\frac{8}{\sqrt[3]{2}}\times \frac{\sqrt[3]{4}}{\sqrt[3]{4}}\] \[=\frac{8\sqrt[3]{4}}{\sqrt[3]{8}}\] \[=\frac{8\sqrt[3]{4}}{2}=4\sqrt[3]{4}\]

OpenStudy (anonymous):

Ah okay. Well then multiply by sqrt(2) / sqrt(2)

OpenStudy (anonymous):

okay thank you two so much. (:

OpenStudy (anonymous):

That gives you (8*sqrt(2)) / (3 * sqrt(2)^2) = (8sqrt(2)) / (3*2) = (8sqrt2) / 6

OpenStudy (anonymous):

My pleasure =)

OpenStudy (anonymous):

@smoothmath why could it not be \[\sqrt[3]{2}\]?

OpenStudy (anonymous):

Um... because I asked her and she said that it's not?

OpenStudy (anonymous):

i am not saying it is or isn't i am just asking how you know

OpenStudy (anonymous):

i certainly don't know, it could be either.

OpenStudy (anonymous):

Meg did you delete your reply? I could have sworn she replied clarifying that is was 3*sqrt(2)

OpenStudy (anonymous):

no it actually was \[\sqrt[3]{2}\]

OpenStudy (anonymous):

@meg it is one of the two answers above. depends on the question really

OpenStudy (anonymous):

oh ok. then it is \[4\sqrt[3]{4}\]

OpenStudy (anonymous):

Haha oh okay. Yes. What satellite did. Not me.

OpenStudy (anonymous):

i accidently said the other one. but i appreciate both of you helping. (: @satellite thanks yours was the answer. an @smooth thank you anyways. lol.

OpenStudy (anonymous):

i have a ton of these problems for homework. but i dont understand the m too well.. @SmoothMath & @satellite73

OpenStudy (anonymous):

post and you will get answers for sure please though remember that \[\sqrt[4]{x}\] is the "fourth root" of x, not 4square root x

OpenStudy (anonymous):

alright thank you ! (:

OpenStudy (anonymous):

Meg, basically the idea is to first "rationalize" the denominator. That's just fancy talk for "get rid of the radical on the bottom."

OpenStudy (anonymous):

Now we know we can multiply the fraction by 1 and not change it's value, so we multiply by some fraction with the same thing on top and bottom.